# Deep Inverse Design for High-Level Synthesis

### v1

@@ V-A Benchmarks
$1$) correlation has three sets of complex loop nests with intensive computation, introducing diversity in the loop structures. Its loop optimization requires bandwidth match from memory access arrangement.
$2$) covariance has two sets of complex loop nests with intensive computation. Its complex data access pattern enlarges room of optimization on parallelization.
$3$) gramSchmidt has one loop nest with five levels performing complicated calculations. The challenges related to both the depth and intricacy of the loop structure include addressing data dependencies and managing loop overhead.
@@
$5$) sort_radix heavily depends on memory access patterns. Optimization on loops, arrays, and function calls improves the efficiency in rearranging the data based on digit positions.
@@ V-B Metric for evaluation
We adopt ADRS for evaluating the methods in our work [3, 4]. A design is said to be non-dominated if there is no other design that improves one objective, e.g., latency and area, without worsening the other, and Pareto front is the set of all such non-dominated designs. ADRS calculates the distance from an approximate Pareto front to the reference Pareto front:
@@
where $f(\gamma,\omega)=max\{|\frac{a_{\omega}-a_{\gamma}}{a_{\gamma}}|,|\frac{l_{\omega}-l_{\gamma}}{l_{\gamma}}|\}$, $\Gamma$ is the reference Pareto front, and $\Omega$ is the approximate Pareto front. For each benchmark, $\Omega$ is obtained from the synthesis results of all the tests conducted by one method or ablation configuration, and we obtain $\Gamma$ from all the $\Omega$ datasets. We do not exhaustively search the design spaces because they are too large. For example, that of stencil_3d contains over $10^{15}$ designs. Smaller ADRS denotes closer distance and better performance.
@@ V-C Comparison with the baselines
The two predictors in DID4444HLS share identical architecture, one for latency (measured in clock cycles) and the other for area (measured in the count of FF, LUT, DSP, BRAM and URAM). The GATv2222 layers are implemented using the PyG library [42].
In our method, we set $\bm{\lambda}=[0,0.2,0.4,0.6,0.8,1]$ to approximate the Pareto front. In each iteration, $5$ designs are sampled for each optimization weight and $30$ designs in total are synthesized. The PIR CDFGs are extracted from the .$*.adb$ files. We choose four state-of-the-art DSE methods for HLS as the baselines for the comparative analysis. All the baselines share the same design sampling process as DID4444HLS for a fair comparison. We introduce the baselines and provide their implementations below:
$1$) GRASP [43] is a tree-based heuristic method composed of two steps: Greedy Randomized Construction (GRC) and Local Search (LS). In each iteration, GRC creates a Restricted Candidate List (RCL) containing the most promising pragmas with minimum costs that are estimated by the predictor based on a decision tree. Then the LS phase further refines the pragmas based on the RCL. All the newly synthesized designs update the predictor for the next iteration. We set $100$ estimators for the random forest regressor, cost distance $\alpha$ as $0.7$, and RCL synthesis interval as $8$. At the beginning, we randomly sample $120$ designs and synthesize them to train the predictor. Then we start the GRASP process in iterations until the synthesis budget is reached.
@@
$3$) SA-ML [16] is an annealing-based method. The predictive models are based on XGBoost [46]. We adopt Clang [^2] https://clang.llvm.org to compile the C++ code with pragmas into LLVM IRs. Then, the design features as the model input are extracted from the IR CDFGs through Programl [20], including attributes such as node counts of each type, path length, and pipelining weight. We set the initial temperature to $10$ with a temperature change rate of $0.1$ and $50$ runs at each temperature. The predictive models are trained with an initial set of $150$ designs randomly sampled and synthesized. Hyperparameter tuning is conducted using Comet.ml with Bayesian optimization [^3] https://www.comet.com. We define a set of $30$ opotimization weights evenly spaced between 00 and $1$ to approximate a Pareto front, and randomly sample $30$ designs as the starting point for each weight. Through simulated annealing, each design transitions to its neighbor by randomly selecting one of its pragma and changing to an adjacent value until the stopping condition, which is set to $500$ rejects or the temperature of $10^{-7}$. The candidates are synthesized after the final transitions.
@@
$4$) GNN-DSE [13] is an exhaustive search method with a GNN surrogate. We adopt Clang to compile C++ code with pragmas into LLVM IRs. Their CDFGs, obtained through Programl, are input into the GAT-based models for predictions. The feature selection mechanism Jumping Knowledge Network (JKN) identifies dominant node features from the GAT layers [35] and maps them into graph features. Subsequently, an MLP follows to output predictions. We construct two separate models to predict latency and area. The models are trained with an initial set of $150$ designs randomly sampled and synthesized. In the DSE phase, we randomly sample $5,000$ designs and predict their latency and area with the trained models. Then, we approximate a Pareto front and randomly select $30$ designs from the front to synthesize.
@@
Tab. I summarizes the performance comparison between DID4444HLS and the baselines. Our method performed the best on all the benchmarks. The average ADRS improvement achieved by DID4444HLS is $42.5\%$, with up to $53.9\%$ for aes. In the worst case (stencil_3d), our method improved the ADRS by $31.1\%$. To further demonstrate the effectiveness of our work, Fig. 6 shows the distances of the approximate Pareto fronts by DID4444HLS and the best-performing baselines to the reference Pareto front. Our method obtained the optimal balance between latency and area for correlation, and these designs overlap the reference Pareto points. Our method achieved more Pareto points than the best-performing baseline for covariance ($16$ vs. $6$), leaving more room for design trade-offs. For aes, our method outperforms the best-performing baseline to such an extent that a gap forms between their respective fronts. Therefore, DID4444HLS demonstrates the superior performance and robustness in improving design QoR.
@@
Method correlation covariance gramSchmidt aes sort_radix stencil_3d GRASP $0.404$ $0.599$ $0.282$ $0.387$ $0.676$ $0.492$ AutoHLS $0.378$ $0.546$ $0.387$ $0.232$ $0.964$ $0.506$ SA-ML $0.639$ $1.148$ $0.199$ $0.335$ $0.307$ $0.498$ GNN-DSE $0.551$ $0.365$ $0.266$ $0.491$ $0.374$ $0.426$ DID4444HLS (ours) $\mathbf{0.199}$ $\mathbf{0.209}$ $\mathbf{0.115}$ $\mathbf{0.107}$ $\mathbf{0.191}$ $\mathbf{0.325}$ Improvement $47.4\%$ $42.7\%$ $42.2\%$ $53.9\%$ $37.8\%$ $31.1\%$
@@
We also compared the time consumed by each method. Fig. 7 lists the average time over all the benchmarks. We set the synthesis timer to five minutes, indicating that the design synthesis process will be terminated if it is not completed within this timeframe. AutoHLS consumed the least time of $30.5$ minutes. SA-ML consumed the most time with $145.5$ minutes. Our method consumed $48.0$ minutes, achieving high efficiency.
@@ V-D Ablation study
GAN consists of a generator and a discriminator. The generator takes random noise as input and generates samples similar to the training data. The discriminator is responsible for distinguishing the generated samples from the real. As training proceeds, both the generator and discriminator improve at their respective tasks. Fig. 8 shows the structure of GAN-based estimators, wherein both the generator and discriminator have their own MLP projector that transforms $\bm{\Theta}_{k}$ into a vector concatenated to the input for learning the conditional distribution, $\mathbf{f}^{\prime}$ denotes the generated graph feature whose ground truth $\mathbf{f}$ is obtained from the extractor of the predictor. After training, we adopt the generator to build the surrogate.
@@
Diffusion model stems from the idea that a continuous Gaussian diffusion process can be reversed with the same functional form as the forward process. The reverse process converges pure noise to samples in the target distribution. The diffusion process can be approximated with a discrete process given enough diffusion steps. The backbone of diffusion model is responsible for extracting meaningful features from the input data at various diffusion steps to iteratively remove noise. Fig. 9 shows the backbone structure we adapted from DiffWave [49], because the gated structure excels at filtering out critical information from the input data, meanwhile skip connections mitigate the vanishing and exploding gradient problems and capture information at multiple levels of abstraction [49, 50, 51]. We set the number of diffusion steps as $50$ with a quadratic noise schedule from $\beta_{1}=10^{-4}$ to $\beta_{50}=0.5$ [50, 51].
@@
Tab. II lists the ADRS values of the ablation configurations, showing the performance to approximate the Pareto fronts achieved by using GAN vs. VAE, diffusion model vs. VAE, and LLVM IR vs. PIR. The complexities of the GAN-based and diffusion model-based estimators are aligned with that of the VAE-based estimators using PyTorch-OpCounter [^4] https://github.com/Lyken17/pytorch-OpCounter, each having around three million parameters. The results show the superior performance of PIR combined with VAE. The average improvement is $48.0\%$ over the best-performing ablation configurations, with up to $58.2\%$ improvement for gramSchmidt. The least improvement is $35.6\%$ for stencil_ 3d.
@@
Method correlation covariance gramSchmidt aes sort_radix stencil_3d PIR + GAN $0.706$ $1.156$ $0.484$ $0.289$ $0.472$ $0.672$ PIR + dffusion model $0.729$ $0.727$ $0.275$ $0.431$ $0.484$ $0.890$ LLVM IR + VAE $0.468$ $0.327$ $0.310$ $0.201$ $0.413$ $0.505$ PIR + VAE (ours) $\mathbf{0.199}$ $\mathbf{0.209}$ $\mathbf{0.115}$ $\mathbf{0.107}$ $\mathbf{0.191}$ $\mathbf{0.325}$ Improvement $57.5\%$ $36.1\%$ $58.2\%$ $46.8\%$ $53.8\%$ $35.6\%$
For each benchmark, we set the synthesis budget to be $180$ solutions for all the methods and ablation configurations. We made modifications to the open-sourced code of GRASP [^5] https://github.com/nibst/GRASP_DSE, and re-implemented the remaining baselines as closely as possible according to their respective descriptions. All the experiments were implemented using Python on the PyTorch framework [41]. The computational hardware utilized included an Intel i7777-$12700$K processor for HLS and an Nvidia RTX $3090$ graphics card for training deep models. We used Vitis HLS as the HLS tool. Our simulation platform was set as the Defense-grade Virtex-$7$Q FPGA with part number xq7777vx690690690690t-rf1930193019301930-$1$I, with a clock constraint of $10ns$. All built-in automatic optimizations were disabled.
@@ VII Conclusion
In this work, we propose DID4444HLS, a method that learns HLS behaviors, enabling design samples to iterate toward the Pareto front with a low synthesis budget. We demonstrate the effectiveness of DID4444HLS with its improvement on design QoR through the tuning of pragmas for loop pipelining, loop unrolling, array partitioning, and function inlining. The comparative experiments exhibit the superior performance of our method over the four state-of-the-art baselines. It achieved an average improvement of $42.5\%$ on ADRS over the best-performing baselines with high robustness and efficiency. In addition, the ablation study validated the construction of DID4444HLS.

### v2

@@ V-A Benchmarks
$1$) correlation has three sets of complex loop nests with intensive computation, introducing diversity in the loop structures. Its loop optimization requires a bandwidth match from memory access arrangement.
$2$) covariance has two sets of complex loop nests with intensive computation. Its complex data access pattern enlarges the room for optimization in parallelization.
$3$) gramSchmidt has one loop nest with five levels performing complicated calculations. The challenges related to both the depth and complexity of the loop structure include addressing data dependencies and managing loop overhead.
@@
$5$) sort_radix heavily depends on memory access patterns. Optimization in loops, arrays, and function calls improves the efficiency of rearranging the data based on digit positions.
@@ V-B Metric for evaluation
We adopt ADRS in evaluating the methods in our work [3, 4]. A design is said to be non-dominated if there is no other design that improves one objective, e.g., latency and resource consumption, without worsening the other, and Pareto front is the set of all such non-dominated designs. ADRS calculates the distance from an approximate Pareto front to the reference Pareto front:
@@
where $f(\gamma,\omega)=max\{|\frac{r_{\omega}-r_{\gamma}}{r_{\gamma}}|,|\frac{l_{\omega}-l_{\gamma}}{l_{\gamma}}|\}$, $\Gamma$ is the reference Pareto front, and $\Omega$ is the approximate Pareto front. For each benchmark, $\Omega$ is obtained from the synthesis results of all the tests conducted by one method or ablation configuration, and we obtain $\Gamma$ from all the $\Omega$ datasets. We do not exhaustively search the design spaces because they are too large. For example, that of stencil_3d contains over $10^{15}$ designs. Smaller ADRS denotes closer distance and better performance.
@@ V-C Comparison with the baselines
The two predictors in DID4444HLS share identical architecture, one for latency measured in clock cycles and the other for resource consumption measured in Sum of Normalized Resource Use (SNRU) [15, 42, 43]:
$SNRU=\frac{1}{4}\cdot(\frac{FF}{FF^{*}}+\frac{LUT}{LUT^{*}}+\frac{DSP}{DSP^{*}}+\frac{BRAM}{BRAM^{*}})$ (9)
where * represents available resource. The GATv2222 layers are implemented using the PyG library [44].
In our method, we set $\bm{\lambda}=[0,0.2,0.4,0.6,0.8,1]$ to approximate the Pareto front. In each iteration, $5$ designs are sampled for each optimization weight and $30$ designs in total are synthesized. $6$ iterations are executed throughout the optimization for each benchmark. The PIR CDFGs are extracted from the .$*.adb$ files. We choose four state-of-the-art DSE methods for HLS as the baselines for the comparative analysis. All the baselines share the same design sampling process as DID4444HLS for a fair comparison. We introduce the baselines and provide their implementations below:
$1$) GRASP [45] is a tree-based heuristic method composed of two steps: Greedy Randomized Construction (GRC) and Local Search (LS). In each iteration, GRC creates a Restricted Candidate List (RCL) containing the most promising pragmas with minimum costs that are estimated by the predictor based on a decision tree. Then the LS phase further refines the pragmas based on the RCL. All the newly synthesized designs update the predictor for the next iteration. We set $100$ estimators for the random forest regressor, cost distance $\alpha$ as $0.7$, and RCL synthesis interval as $8$. In the beginning, we randomly sample $120$ designs and synthesize them to train the predictor. Then we start the GRASP process in iterations until the synthesis budget is reached.
@@
$3$) SA-ML [16] is an annealing-based method. The predictive models are based on XGBoost [48]. We adopt Clang [^2] https://clang.llvm.org to compile the C++ code with pragmas into LLVM IRs. Then, the design features as input to the model are extracted from the IR CDFGs through Programl [20], including attributes such as the number of nodes of each type, the path length, and the pipelining weight. We set the initial temperature to $10$ with a temperature change rate of $0.1$ and $50$ runs at each temperature. Predictive models are trained with an initial set of $150$ designs randomly sampled and synthesized. Hyperparameter tuning is performed using Comet.ml with Bayesian optimization [^3] https://www.comet.com. We define a set of $30$ optimization weights evenly spaced between 00 and $1$ to approximate a Pareto front, and randomly sample $30$ designs as the starting point for each weight. Through simulated annealing, each design transitions to its neighbor by randomly selecting one of its pragma and changing to an adjacent value until the stopping condition, which is set to $500$ rejects or the temperature of $10^{-7}$. The candidates are synthesized after the final transitions.
@@
$4$) GNN-DSE [13] is an exhaustive search method with a GNN surrogate. We adopt Clang to compile C++ code with pragmas into LLVM IRs. Their CDFGs, obtained through Programl, are input into the GAT-based models for predictions. The Jumping Knowledge Network (JKN) feature identifies the dominant node features of the GAT layers [35] and maps them to graph features. Subsequently, an MLP follows to output predictions. We construct two separate models to predict latency and resource consumption. The models are trained with an initial set of $150$ designs randomly sampled and synthesized. In the DSE phase, we randomly sample $10,000$ designs for a comparable runtime and predict their latency and resource consumption with trained models. Then, we approximate a Pareto front and randomly select $30$ designs from the front to synthesize.
@@
Tab. I summarizes the performance comparison between DID4444HLS and the baselines. Our method performed the best on all the benchmarks. The average ADRS improvement over the best-performing baselines achieved by DID4444HLS is $42.8\%$, with up to $74.3\%$ for sort_radix. In the worst case (aes), our method improved the ADRS by $23.1\%$. To further demonstrate the effectiveness of our work, Fig. 6 shows the distances of the approximate Pareto fronts by DID4444HLS and the best-performing baselines to the reference Pareto front. Our method offers a wider spread of solutions in latency values for correlation, making it more versatile for a range of design requirements. For aes, our method significantly outperforms the best-performing baseline, resulting in a distinct gap between the two fronts. Our method achieved more Pareto points than the best-performing baseline for stencil_3d ($8$ vs. $1$), uncovering broader trade-offs and revealing greater potential for diverse designs. Therefore, DID4444HLS demonstrates the superior performance and robustness in improving the design QoR.
@@
Method correlation covariance gramSchmidt aes sort_radix stencil_3d GRASP $0.262$ $0.481$ $0.359$ $0.515$ $0.701$ $0.612$ AutoHLS $0.402$ $0.473$ $0.270$ $0.286$ $1.033$ $0.597$ SA-ML $0.780$ $0.766$ $0.170$ $0.286$ $0.514$ $0.487$ GNN-DSE $0.356$ $0.486$ $0.330$ $0.356$ $0.610$ $0.436$ DID4444HLS (ours) $\mathbf{0.191}$ $\mathbf{0.166}$ $\mathbf{0.107}$ $\mathbf{0.220}$ $\mathbf{0.132}$ $\mathbf{0.305}$ Improvement $27.1\%$ $64.9\%$ $37.1\%$ $23.1\%$ $74.3\%$ $30.0\%$
@@
We also compared the time consumed by each method. Fig. 7 lists the average time over all the benchmarks. Since designs can take forever to synthesize, we set the synthesis timer to ten minutes, indicating that the design synthesis process will be terminated if it is not completed within this timeframe. AutoHLS consumed the least time of $49.4$ minutes. SA-ML consumed the most time with $185.6$ minutes. Our method consumed $76.4$ minutes. Although not the fastest, the runtime of our method is justified given its superior performance.
@@ V-D Ablation study
GAN consists of a generator and a discriminator. The generator takes random noise as input and generates samples similar to the training data. The discriminator is responsible for distinguishing the generated samples from the real ones. As training progresses, both the generator and the discriminator improve in their respective tasks. Fig. 8 shows the structure of GAN-based estimators, where both the generator and the discriminator have their own MLP projector that transforms $\bm{\Theta}_{k}$ into a vector concatenated to the input for learning the conditional distribution, $\mathbf{f}^{\prime}$ denotes the generated graph feature whose ground truth $\mathbf{f}$ is obtained from the predictor’s extractor. After training, we adopt the generator to build the surrogate.
@@
Diffusion model stems from the idea that a continuous Gaussian diffusion process can be reversed with the same functional form as the forward process. The reverse process converges pure noise to samples in the target distribution. The diffusion process can be approximated with a discrete process given enough diffusion steps. The backbone of diffusion model is responsible for extracting meaningful features from the input data at various diffusion steps to iteratively remove noise. Fig. 9 shows the backbone structure we adapted from DiffWave [51], because the gated structure excels at filtering out critical information from the input data. Meanwhile, skip connections mitigate vanishing and exploding gradient problems and capture information at multiple levels of abstraction [51, 52, 53]. We set the number of diffusion steps as $50$ with a quadratic noise schedule from $\beta_{1}=10^{-4}$ to $\beta_{50}=0.5$ [52, 53].
@@
Tab. II lists the ADRS values of the ablation configurations, showing the performance to approximate the Pareto fronts achieved by using GAN vs. VAE, diffusion model vs. VAE, and LLVM IR vs. PIR. The complexities of the GAN-based and diffusion model-based estimators are aligned with that of the VAE-based estimators using PyTorch-OpCounter [^4] https://github.com/Lyken17/pytorch-OpCounter, each having around three million parameters. The results show the superior performance of PIR combined with VAE. The average improvement is $43.0\%$ over the best-performing ablation configurations, with up to $71.1\%$ improvement for sort_radix. The least improvement is $19.1\%$ for aes.
@@
Method correlation covariance gramSchmidt aes sort_radix stencil_3d PIR + GAN $0.405$ $0.711$ $0.569$ $0.340$ $0.456$ $0.601$ PIR + dffusion model $0.292$ $0.659$ $0.506$ $0.407$ $1.065$ $1.053$ LLVM IR + VAE $0.448$ $0.294$ $0.208$ $0.272$ $0.514$ $0.518$ PIR + VAE (ours) $\mathbf{0.191}$ $\mathbf{0.166}$ $\mathbf{0.107}$ $\mathbf{0.220}$ $\mathbf{0.132}$ $\mathbf{0.305}$ Improvement $34.6\%$ $43.5\%$ $48.6\%$ $19.1\%$ $71.1\%$ $41.1\%$
For each benchmark, we set the synthesis budget to be $180$ solutions. Thus, each method or ablation configuration is allowed to synthesize at most $180$ designs to approximate the Pareto front. We made modifications to the open-sourced code of GRASP [^5] https://github.com/nibst/GRASP_DSE, and re-implemented the remaining baselines as closely as possible according to their respective descriptions. All the experiments were implemented using Python on the PyTorch framework [41]. The computational hardware utilized included an Intel i7777-$12700$K processor for HLS and an Nvidia RTX $3090$ graphics card for training deep models. We used Vitis HLS $2023.1$ as the HLS tool. Our simulation platform was set as the Defense-grade Virtex-$7$Q FPGA with part number xq7777vx690690690690t-rf1930193019301930-$1$I, with a clock constraint of $10ns$. All built-in automatic optimizations were disabled.
@@ VII Conclusion
In this work, we propose DID4444HLS, a method that learns HLS behaviors, enabling design samples to iterate toward the Pareto front with a low synthesis budget. We demonstrate the effectiveness of DID4444HLS with its improvement on design QoR through the tuning of pragmas for loop pipelining, loop unrolling, array partitioning, and function inlining. The comparative experiments exhibit the superior performance of our method over the four state-of-the-art baselines. It achieved an average improvement of $42.8\%$ on ADRS over the best-performing baselines with high robustness and efficiency. In addition, the ablation study validated the construction of DID4444HLS.